OpenStudy (dehelloo):

Susan can throw a hockey ball so that its path is given by the equation What is the highest level the hockey ball reaches. The answer is 16 but how do you get that from h=-x(x-8)

OpenStudy (anonymous):

If you distribute, you can turn it into a quadratic. \[h=-x(x-8)\rightarrow h=-x^2-8x\]we set it to zero\[-x^2-8x+0=0\] Now we can use the following equation to check for our max/min value. \[c-\frac{ b^2 }{ 4a }\rightarrow0-\frac{ (-8)^2 }{ 4(-1) }\rightarrow \frac{ -64 }{ -4 }\rightarrow16\]So we get that the maximum of this prabola is 16, which is the highest the ball will go.